Presentasjon lastes. Vennligst vent

Presentasjon lastes. Vennligst vent

ALICE 2015-16 YETS Arturo Tauro. Outline YETS planning meeting 11/12/152 YETS duration EN-HE, EN-EL, EN-CV, PH-DT (gas) and GS-ASE activities Acoustic.

Liknende presentasjoner


Presentasjon om: "ALICE 2015-16 YETS Arturo Tauro. Outline YETS planning meeting 11/12/152 YETS duration EN-HE, EN-EL, EN-CV, PH-DT (gas) and GS-ASE activities Acoustic."— Utskrift av presentasjonen:

1 ALICE YETS Arturo Tauro

2 Outline YETS planning meeting 11/12/152 YETS duration EN-HE, EN-EL, EN-CV, PH-DT (gas) and GS-ASE activities Acoustic noise measurements (part-II) Detector activities Wiener LVPS consolidation, Dipole consolidation CERN closure: w52 and w53 Gas and cooling status P2 ‘ronde’ and CERN access Visits

3 Remarks YETS planning meeting 11/12/153 L3 doors stay closed Beampipe stays under vacuum  MNF restricted access Access mode in general Access starts on Mon 14 th Dec at 9:00 AM RP check Andre + Philippe/Klaus SL to ramp down magnets at 6:00 AM

4 YETS duration YETS planning meeting 11/12/154 Year End Technical Stop duration: 14 th Dec 2015 – 6 th March weeks (including 2 Xmas break) NB: w1 starts on Jan 4 th ! Décembre 2015 Janvier 2016 Février Mars 4949 mar1 ven1 Jour de l'An 5 lun1 9 mar1 mer2 sam2 mar2 mer2 jeu3 dim3 mer3 jeu3 ven4 1 lun4 jeu4 ven4 sam5 mar5 ven5 sam5 dim6 mer6 sam6 dim lun7 jeu7 dim lun7 mar8 ven8 6 lun8 mar8 mer9 sam9 mar9 AUG mer9 jeu10 dim10 mer10 jeu10 ven11 2 lun11 jeu11 ven11 sam12 mar12 ven12 sam12 dim13 mer13 sam13 dim lun14 jeu14 dim14 1 lun14 mar15 ven15 7 lun15 mar15 mer16 sam16 mar16 mer16 jeu17 dim17 mer17 jeu17 ven18 3 lun18 jeu18 ven18 sam19 mar19 ven19 sam19 dim20 mer20 sam20 dim lun21 jeu21 dim lun21 mar22 ven22 8 lun22 mar22 mer23 sam23 mar23 mer23 jeu24 Noël dim24 mer24 jeu24 ven25 4 lun25 jeu25 ven25 Vendredi saint sam26 mar26 ven26 DSO sam26 dim27 mer27 sam27 dim lun28 jeu28 dim lun28 Lundi de Pâques mar29 ven29 lun29 mar29 mer30 sam30 mer30 jeu31 dim31 jeu31

5 EN-HE: replacement cavern lift YETS planning meeting 11/12/155 Décembre 2015 Janvier 2016 Février Mars 4949 mar1 ven1 Jour de l'An 5 lun1 9 mar1 mer2 sam2 mar2 mer2 jeu3 dim3 mer3 jeu3 ven4 1 lun4 jeu4 ven4 sam5 mar5 ven5 sam5 dim6 mer6 sam6 dim lun7 jeu7 dim lun7 mar8 ven8 6 lun8 mar8 mer9 sam9 mar9 mer9 jeu10 dim10 mer10 jeu10 ven11 2 lun11 jeu11 ven11 sam12 mar12 ven12 sam12 dim13 mer13 sam13 dim lun14 jeu14 dim14 1 lun14 mar15 ven15 7 lun15 mar15 mer16 sam16 mar16 mer16 jeu17 dim17 mer17 jeu17 ven18 3 lun18 jeu18 ven18 sam19 mar19 ven19 sam19 dim20 mer20 sam20 dim lun21 jeu21 dim lun21 mar22 ven22 8 lun22 mar22 mer23 sam23 mar23 mer23 jeu24 Noël dim24 mer24 jeu24 ven25 4 lun25 jeu25 ven25 Vendredi saint sam26 mar26 ven26 sam26 dim27 mer27 sam27 dim lun28 jeu28 dim lun28 Lundi de Pâques mar29 ven29 lun29 mar29 mer30 sam30 mer30 jeu31 dim31 jeu31 2 shifts, work also during Xmas closure HSE reception 29 Jan Partial opening PX24 plug (w.51) ALICE personnel via PM25 (no need self-rescue masks) If PM25 lift out of order, no need to evacuate

6 EN-HE YETS planning meeting 11/12/156 Reinstall UX crane electronics on Mon Partial opening PX24 plug (w.51) No access UX25 cavern on Tue (exceptions: TPC, SSD) Décembre 2015 Janvier 2016 Février Mars 4949 mar1 ven1 Jour de l'An 5 lun1 9 mar1 mer2 sam2 mar2 mer2 jeu3 dim3 mer3 jeu3 ven4 1 lun4 jeu4 ven4 sam5 mar5 ven5 sam5 dim6 mer6 sam6 dim lun7 jeu7 dim lun7 mar8 ven8 6 lun8 mar8 mer9 sam9 mar9 AUG mer9 jeu10 dim10 mer10 jeu10 ven11 2 lun11 jeu11 ven11 sam12 mar12 ven12 sam12 dim13 mer13 sam13 dim lun14 jeu14 dim14 1 lun14 mar15 ven15 7 lun15 mar15 mer16 sam16 mar16 mer16 jeu17 dim17 mer17 jeu17 ven18 3 lun18 jeu18 ven18 sam19 mar19 ven19 sam19 dim20 mer20 sam20 dim lun21 jeu21 dim lun21 mar22 ven22 8 lun22 mar22 mer23 sam23 mar23 mer23 jeu24 Noël dim24 mer24 jeu24 ven25 4 lun25 jeu25 ven25 Vendredi saint sam26 mar26 ven26 DSO sam26 dim27 mer27 sam27 dim lun28 jeu28 dim lun28 Lundi de Pâques mar29 ven29 lun29 mar29 mer30 sam30 mer30 jeu31 dim31 jeu :00 AM to 8:00 AM: maintenance UX lift W52/53: SXL2 crane inspection (Bonfanti) W1: maintenance UX crane

7 EN-EL: test secours https://edms.cern.ch/ui/file/ /1/ENNC_EL_2015_258.pdf YETS planning meeting 11/12/157 Planned for next Thursday, 17 th Dec, at 6:00 AM (10’) For P2 this will affect: all buildings including SX2 and SXL2 all offices, meeting rooms, wall outlets, light, coffee machine the non UPS part of the ARC the non UPS part of the detector cooling plants (interlocks triggered) the gas systems (PLCs stopped, interlocks triggered)  Albin to stop gas systems beforehand the circulation of iced and mixed water, ventilation (will restart after the 10 min. interruption) Are NOT affected at P2: all equipment on UPS (notably all CRs and inside ARC) all racks in CR and UX As we do not know if all ventilation will restart OK, we will suspend cavern access until we confirm that all is nominal

8 EN-EL: other tests YETS planning meeting 11/12/158 Test auto transfert: 30 th Jan In principle transparent for ALICE, but does affect Meyrin AUG test: 9 th Feb - no access P2 Cut all, no access underground Maintenance on 2 transformers: EMT408/25 and EMT307/2S, foreseen right after the AUG test, on 10 th Feb Cut all racks in cavern (EXD) Cut SXL2 & cleanroom Other activities: Put ITS ventilator under UPS Electrical distribution SXL2 cleanroom (w4) Install light L3 power converter (SX2)

9 EN-CV: maintenance primary cooling and ventilation YETS planning meeting 11/12/159 Stop cooling towers: w5-6-7 Stop chilled-, mixed- and demineralized-water: w.6 – NO cooling P2. Install magnetic filter UW25  stop CW circulation (1 day), right after the AUG Ventilation: -L3 w51 (several 3’ interruptions) -UX25 w5 -SG w5 or w6 – no flammable gases for 3 days -MCH/MTG w7 -L3 w9 – no access inside L3 on Mon 29 th Feb Décembre 2015 Janvier 2016 Février Mars 4949 mar1 ven1 Jour de l'An 5 lun1 9 mar1 mer2 sam2 mar2 mer2 jeu3 dim3 mer3 jeu3 ven4 1 lun4 jeu4 ven4 sam5 mar5 ven5 sam5 dim6 mer6 sam6 dim lun7 jeu7 dim lun7 mar8 ven8 6 lun8 mar8 mer9 sam9 mar9 AUG mer9 jeu10 dim10 mer10 jeu10 ven11 2 lun11 jeu11 ven11 sam12 mar12 ven12 sam12 dim13 mer13 sam13 dim lun14 jeu14 dim14 1 lun14 mar15 ven15 7 lun15 mar15 mer16 sam16 mar16 mer16 jeu17 dim17 mer17 jeu17 ven18 3 lun18 jeu18 ven18 sam19 mar19 ven19 sam19 dim20 mer20 sam20 dim lun21 jeu21 dim lun21 mar22 ven22 8 lun22 mar22 mer23 sam23 mar23 mer23 jeu24 Noël dim24 mer24 jeu24 ven25 4 lun25 jeu25 ven25 Vendredi saint sam26 mar26 ven26 DSO sam26 dim27 mer27 sam27 dim lun28 jeu28 dim lun28 Lundi de Pâques mar29 ven29 lun29 mar29 mer30 sam30 mer30 jeu31 dim31 jeu31

10 EN-CV: detector cooling maintenance YETS planning meeting 11/12/1510 PlantDays of cutCommentOther remarks SSD/SDD (filter included) 10/2 and 11/2No cooling during this period Could have some perturbations on the temp due to primary circuits maintenance from week 5 till end of 7 TPC FEC1/2 and 2/2No cooling during this period HMPID12/2 and 15/2No cooling during this period SPD15/2 to 17/2No cooling during this period TRD3/2 and 4/2No cooling during this period TOF PHOS EMC DCA4/2 and 5/2No cooling during this period TPC ROD1/2 and 2/2No cooling during this period Replacement of the batteries on the communication PLC 9/2 No com during 1 hour, it will be done after the AUG tests RP samples will be collected on Mon 14 th Dec 9:00 AM (including Dipole, Solenoid, PHOS C6F14) Smear tests and collection of SAMs  Klaus?

11 EN-CV: new ITS dryer YETS planning meeting 11/12/1511 Replacement unit in SX2 + duct Modification N2 supply (fire- extinguisher, circuit ITS) Planning: -Switchover to FISAIR (CR5) in w51 + test (Ton) -Switch back to Dessica end w51 -Dismantling of Dessica starts in w1 -New pipe: surface part installed w49, cavern part in w3-4 -Commissioning: w5-6 Décembre 2015 Janvier 2016 Février Mars 4949 mar1 ven1 Jour de l'An 5 lun1 9 mar1 mer2 sam2 mar2 mer2 jeu3 dim3 mer3 jeu3 ven4 1 lun4 jeu4 ven4 sam5 mar5 ven5 sam5 dim6 mer6 sam6 dim lun7 jeu7 dim lun7 mar8 ven8 6 lun8 mar8 mer9 sam9 mar9 AUG mer9 jeu10 dim10 mer10 jeu10 ven11 2 lun11 jeu11 ven11 sam12 mar12 ven12 sam12 dim13 mer13 sam13 dim lun14 jeu14 dim14 1 lun14 mar15 ven15 7 lun15 mar15 mer16 sam16 mar16 mer16 jeu17 dim17 mer17 jeu17 ven18 3 lun18 jeu18 ven18 sam19 mar19 ven19 sam19 dim20 mer20 sam20 dim lun21 jeu21 dim lun21 mar22 ven22 8 lun22 mar22 mer23 sam23 mar23 mer23 jeu24 Noël dim24 mer24 jeu24 ven25 4 lun25 jeu25 ven25 Vendredi saint sam26 mar26 ven26 DSO sam26 dim27 mer27 sam27 dim lun28 jeu28 dim lun28 Lundi de Pâques mar29 ven29 lun29 mar29 mer30 sam30 mer30 jeu31 dim31 jeu31

12 BE-CO control infrastructure maintenance YETS planning meeting 11/12/1512 A major upgrade is planned by BE-CO for the first week of January – specifically Mon 4 th until Wed 6 th inclusive – which will have a significant effect on access to the WinccOA (PVSS) applications/remote desktop etc for the gas systems.

13 PH-DT and EN-MEF: gas activities YETS planning meeting 11/12/1513 Stop flammable gases (HMP and MTG) on Monday Dec 14 th (Ar-H2 will stay over Xmas closure). Purge systems (Albin) Stop N2 dewar on Mon Dec 14 th at 9:00 (full-day) (connection N2 transfer line)  switch to batteries (480m3 batteries on site, enough for at least 24h operation. Consumption 6-7m3/h. David to buy more batteries). N2 consumption higher (+1m3/h) on Wed-Thu (w51) + w3- 4 (tbc) (SSD N2 injection). EN-MEF (David): N2 dewar will be filled up before CERN closure PH-DT: remove safety valves TRD gas system. Document to be issued by Roberto PH-DT (Albin): recuperation and compression of TRD Xe into the B50 gas cylinder during the two weeks of CERN closure

14 GS-ASE YETS planning meeting 11/12/1514 Safety test on 24-Dec (whole day) – NO ACCESS UX25 Maintenance gas detection (SG2, PX24 and UX25): 17 and 18 Dec. Fri 18 th between 14:00 and 15:00: Several gas systems will be stopped (TPC, TOF, HMP, MTR). SG2 general services will be cut, EL needed to rearm the main breaker Ar-H2 will be stopped SNIFFER maintenance (end Jan, date tbc)  cleaning tubes ALICE responsibility (Didier) Maintenance red phones: 12 Jan Maintenance fire detection (SFDIN LHC2 RACKS): 22 Jan Maintenance doors, SAS etc: 3 to 5 Feb

15 Other activities YETS planning meeting 11/12/1515 Didier: installation fences UX25 visitor’s path INEO: access trap L3 polarity switch (SX2) TE-VSC: close manual valve on Mon 14 th Dec IT: IP renumbering of the ALICE bypass services, early morning Thu 17 Dec AM: It might stop DCS and DAQ for the time of the intervention (1 hour) IT: installation Ethernet outlets inside L3 – w8 or 9 (tbc) DGS-RP: maintenance RP equipment on Jan-11 th : RAMSES Software upgrade to all electronics. Functionality source test for all monitors (chambers, PCM & HFM)

16 Acoustic noise measurements (part-II) YETS planning meeting 11/12/1516 Planned beginning of w.5, one night starting at 22:00 PM Start by switching off all that was already tested during TS2 (ALICE magnets, outside chillers, cavern ventilation,...)  no access underground during the cavern ventilation stop (essentially for full duration of test) The components which were not switched off during TS2 will be switched off in sequence: 1.Cryo 2.Cooling towers 3.Power converters Stop of cooling towers will stop cooling of our racks. CR racks can be switched to 'eau brute', so no issue. Cavern racks will loose cooling, but mixed water can continue to circulate (but will not be cooled), so temperature will increase but if we limit the load we can survive a few hours Switch back on everything in inverse sequence

17 Summary all perturbations red=implies no access to certain zones, orange=limited cooling YETS planning meeting 11/12/1517 Décembre 2015 Janvier 2016 Février Mars 4949 mar1 ven1 Jour de l'An 5 lun1 9 mar1 mer2 sam2 mar2 mer2 jeu3 dim3 mer3 jeu3 ven4 1 lun4 jeu4 ven4 sam5 mar5 ven5 sam5 dim6 mer6 sam6 dim lun7 jeu7 dim lun7 mar8 ven8 6 lun8 mar8 mer9 sam9 mar9 AUG mer9 jeu10 dim10 mer10 jeu10 ven11 2 lun11 jeu11 ven11 sam12 mar12 ven12 sam12 dim13 mer13 sam13 dim lun14 jeu14 dim14 1 lun14 mar15 ven15 7 lun15 mar15 mer16 sam16 mar16 mer16 jeu17 dim17 mer17 jeu17 ven18 3 lun18 jeu18 ven18 sam19 mar19 ven19 sam19 dim20 mer20 sam20 dim lun21 jeu21 dim lun21 mar22 ven22 8 lun22 mar22 mer23 sam23 mar23 mer23 jeu24 Noël dim24 mer24 jeu24 ven25 4 lun25 jeu25 ven25 Vendredi saint sam26 mar26 ven26 DSO sam26 dim27 mer27 sam27 dim lun28 jeu28 dim lun28 Lundi de Pâques mar29 ven29 lun29 mar29 mer30 sam30 mer30 jeu31 dim31 jeu Jan: annual maintenance on the BE- CO controls infrastructure 14 Dec: stop N2 dewar (switch to batteries); 1h DCS interruption 15 Dec: open PX24 shielding (no access UX25) 17 Dec: test secours (6:00AM); IP renumbering of the ALICE bypass services (might stop DAQ, DCS 1hr); ramp-up L3 (17:00-20:00) 18 Dec: gas detection maintenance (cut Ar-H2 and other gas systems at 14:00, for 1h) 24 Dec: safety test (no access UX25) W5-6-7: primary cooling maintenance (reduced cooling) W5 (exact date tbc): acoustic noise measurements (Part-II) (no access UX25 one night after 22:00) 9 Feb: AUG test (no access P2) 10 Feb: maintenance transformers: cut all racks in UX25 26 Feb: DSO test: no access UX25 29 Feb and 1 Mar: maintenance L3 ventilation: no access inside L3 30 Jan: test auto- transfert

18 Detector activities YETS planning meeting 11/12/1518

19 TPC RCU2 YETS planning meeting 11/12/1519 Plan (see next slide): Install 1 sector in w51 Install other sectors w1-3 Test w4-5 Extra manpower needed: Samuel all the time Yannick occasionally (e.g. platforms) PH-DT FSU, 4 days in w3  this person could be available for two weeks: w3 and 4 Décembre 2015 Janvier 2016 Février Mars 4949 mar1 ven1 Jour de l'An 5 lun1 9 mar1 mer2 sam2 mar2 mer2 jeu3 dim3 mer3 jeu3 ven4 1 lun4 jeu4 ven4 sam5 mar5 ven5 sam5 dim6 mer6 sam6 dim lun7 jeu7 dim lun7 mar8 ven8 6 lun8 mar8 mer9 sam9 mar9 AUG mer9 jeu10 dim10 mer10 jeu10 ven11 2 lun11 jeu11 ven11 sam12 mar12 ven12 sam12 dim13 mer13 sam13 dim lun14 jeu14 dim14 1 lun14 mar15 ven15 7 lun15 mar15 mer16 sam16 mar16 mer16 jeu17 dim17 mer17 jeu17 ven18 3 lun18 jeu18 ven18 sam19 mar19 ven19 sam19 dim20 mer20 sam20 dim lun21 jeu21 dim lun21 mar22 ven22 8 lun22 mar22 mer23 sam23 mar23 mer23 jeu24 Noël dim24 mer24 jeu24 ven25 4 lun25 jeu25 ven25 Vendredi saint sam26 mar26 ven26 DSO sam26 dim27 mer27 sam27 dim lun28 jeu28 dim lun28 Lundi de Pâques mar29 ven29 lun29 mar29 mer30 sam30 mer30 jeu31 dim31 jeu31 Installation full TPC Test

20 TPC RCU2 YETS planning meeting 11/12/1520 W51: installation in one sector (C-side) Install RCU2 in one sector on C-side, on Mon 14th Dec Yannick+Samuel: install access structures on the A-side Philippe+Didier: cut bars baby-frame + install table + light Ramp up L3 on Thu from 17:00 PM to 20:00 PM DAQ needed until Friday W52-53: validate RCU2 W1-3: installation in all other sectors — 2 teams of 2 people each (one on A- and one on C-side). Watcher does cooling modification W1: modify 5 sectors on upper A-side and 5 sectors on C-side W2: modify 5 sectors on upper A-side and 6 sectors on C-side W3: modify 5 sectors on A-side, 3 sectors on lower A-side and 7 sectors on C-side – SSD & SPD & FMD boxes removed W4-5: test full TPC with RCU2 (plus fixing remaining problems) Friday in week 5: unmount LV PS (for modification) W6: Wiener LV PS modification (no cooling available) W7: re-install Wieners

21 TPC laser and FMD YETS planning meeting 11/12/1521 TPC laser: one mirror needs probably to be exchanged Short access on Thursday 17 Dec to assess possible damages to the mirrors Then probably one day access sometime during the shutdown, preferably w4 or later, in order to repair Another day needed towards the end of the period for re-alignment FMD: not yet decided whether to change RCU to RCU2. In case, it will take 1-2 days, with access to both A and C side  wait TS1 Remove the FMD RCU board on the A-side (for the TPC work). It will be done either on Dec 17 th or beginning Jan

22 ITS YETS planning meeting 11/12/1522 W51 (SSD, Ton): Mon-Tue: test humidity of SSD with Fisair machine Wed-Thu: installing N2 distribution lines (PP4) and stability test with N2 Fri: disconnect N2 supply and leave SSD running with Dessica for the Xmas break W3: (SSD, Ton): remove PP1B for TPC work on A1 side. After two days the PP1B can be restored (SPD, Michel): disconnect SPD fibers W5: replace one SSD P regulator

23 TRD – w51 YETS planning meeting 11/12/1523 Inspect SM 9 and 17 with Yannick (interference with TPC transfer rails) – either Wed or Thu PM, check with Yannick Try avoiding to cool down the SMs when no power to avoid large Xenon pressure variations (e.g. gas injection). See https://indico.cern.ch/event/462525/. Possible solution: increase the T set-point (cooling plant) when detector is switched-off. Test on Mon 14 th Dec AM. If successful, EN-CV to install new Temp sensor on the return line + modify PLC logic (mid Jan)https://indico.cern.ch/event/462525/ Tue and Wed: removal of the HV current-spike detection system in CR4 together with inspection/connection of optical fibers in the CTP area Nothing planned inside the L3 magnet

24 TOF YETS planning meeting 11/12/1524 This is the list of our planned activities, some of them can be carried out in parallel. All of them are inside L3 (except the ethernet switches). For the HV intervention, technicians from Bologna are needed; we scheduled 1 week in January and 1 week in February (tbc). Replacement of few Local Trigger Modules (LTM) on the TOF crates inside L3 (5 – 9 boards, to be checked)  1 week Few HV connectors to be reworked inside L3. Welding needed: smoke detection off?  1 week Test of new HV distribution boxes  1 week few cooling tubes have been found crushed and bent, some look seriously damaged; replacement? Ethernet redundancy (connect power supply)  in collaboration with DCS

25 PHOS YETS planning meeting 11/12/1525 Maintenance of LED system Work from 15th of December till 1st of January, also during CERN closure  NOT inside L3 Top-up C6F14 (33 litres) for the PHOS matrix cooling plant  towards the end of the YETS SPIE: install rigid air ducts to cool down PHOS RCUs – w7

26 EMC/DCA YETS planning meeting 11/12/1526 Install cooling tubes for the remaining 19 SRUs of EMCAL/DCAL Production ongoing Installation inside L3 in February Remove a few covers in w51 (after Monday)

27 T0A and FIT YETS planning meeting 11/12/1527 T0A: replace short-optical fibers for laser system FIT: install one of the FIT module near T0A (tbd)

28 V0 PMT replacement YETS planning meeting 11/12/1528 One channel in V0A (Ring1,Sector6) is almost dead  need to replace the PMT Proposal: Mon 8 th Feb? Fine for Borge.

29 MCH and MTR Plan still under definition YETS planning meeting 11/12/1529 MCH: Mon 14: voltage measurement in ST1 (doubt about the correctness of the sectors position in DCS wrt the physical sector position) Change few slats Change quadrant of station 2 MTR: Installation/commissioning new regional cards, most probably w4 Maintenance on the gas connectors and on the front-end electronics (not scheduled yet) Replacement of one or more RPCs

30 DCS YETS planning meeting 11/12/1530 DCS we will need max. 1.5 hrs on Monday morning Dec 14 th to make a reference 'snapshot' of the system. MCH and TPC would like to continue operation as much as possible: we will handle them with priority. All other interventions that might seriously disturb operations will be planned in the shadow of the cooling intervention (notably week 6, around the AUG test).

31 AD YETS planning meeting 11/12/1531 Access to both sides (check signals, possibly replace PMTs)

32 ZDC YETS planning meeting 11/12/1532 Replace one PMT (Paolo checks access with J.Coupard) W2-3: stand-alone tests from ARC (DAQ-DCS needed)

33 Wiener LVPS consolidation YETS planning meeting 11/12/1533

34 Consolidation Wiener LVPS YETS planning meeting 11/12/ in total

35 Consolidation Wiener LVPS YETS planning meeting 11/12/1535 Draining ALICE P2 team (Philippe) De-installation & reinstallation detector responsibility Transport ALICE Repair in surface: Philippe + Evgeny + 2 SPIE technicians Clean rack during rework (Philippe) Pressure test before reinstallation Décembre 2015 Janvier 2016 Février Mars 4949 mar1 ven1 Jour de l'An 5 lun1 9 mar1 mer2 sam2 mar2 mer2 jeu3 dim3 mer3 jeu3 ven4 1 lun4 jeu4 ven4 sam5 mar5 ven5 sam5 dim6 mer6 sam6 dim lun7 jeu7 dim lun7 mar8 ven8 6 lun8 mar8 mer9 sam9 mar9 AUG mer9 jeu10 dim10 mer10 jeu10 ven11 2 lun11 jeu11 ven11 sam12 mar12 ven12 sam12 dim13 mer13 sam13 dim lun14 jeu14 dim14 1 lun14 mar15 ven15 7 lun15 mar15 mer16 sam16 mar16 mer16 jeu17 dim17 mer17 jeu17 ven18 3 lun18 jeu18 ven18 sam19 mar19 ven19 sam19 dim20 mer20 sam20 dim lun21 jeu21 dim lun21 mar22 ven22 8 lun22 mar22 mer23 sam23 mar23 mer23 jeu24 Noël dim24 mer24 jeu24 ven25 4 lun25 jeu25 ven25 Vendredi saint sam26 mar26 ven26 DSO sam26 dim27 mer27 sam27 dim lun28 jeu28 dim lun28 Lundi de Pâques mar29 ven29 lun29 mar29 mer30 sam30 mer30 jeu31 dim31 jeu31

36 Consolidation Wiener LVPS YETS planning meeting 11/12/1536 Status Dec.8 th (Andre) All detectors are asked to bring their spares to P2! Total in DBExcluded Total to be done DeinstallReworkReinstallDone Still to be done UXP2Other AD2 2w5w6w7 211 CTP11 w EMC8 8w5w6w HMPID2 2-w MCH40337w4w5w MTR10 w5w6w PHOS10 w4w5w PMD SPD-PIT2 2w3w4w SSD10 w T02 2w3w4w5 22 TOF3 3w3w4w TPC23 w5 (Fri)w6w TRD102 w2w3-4w V0312w3w4w5 211 ALICE

37 Dipole consolidation YETS planning meeting 11/12/1537

38 Dipole cooling repair YETS planning meeting 11/12/1538

39 Dipole cooling repair YETS planning meeting 11/12/1539 Team: Greg + Evgeny + Philippe (w1 only) Schedule: w1 and w2 (w3?) Install drainage line w51 (Philippe, Didier, INEO) + purge cooling Verification solution in w51 Use scaffolding to access connections Décembre 2015 Janvier 2016 Février Mars 4949 mar1 ven1 Jour de l'An 5 lun1 9 mar1 mer2 sam2 mar2 mer2 jeu3 dim3 mer3 jeu3 ven4 1 lun4 jeu4 ven4 sam5 mar5 ven5 sam5 dim6 mer6 sam6 dim lun7 jeu7 dim lun7 mar8 ven8 6 lun8 mar8 mer9 sam9 mar9 AUG mer9 jeu10 dim10 mer10 jeu10 ven11 2 lun11 jeu11 ven11 sam12 mar12 ven12 sam12 dim13 mer13 sam13 dim lun14 jeu14 dim14 1 lun14 mar15 ven15 7 lun15 mar15 mer16 sam16 mar16 mer16 jeu17 dim17 mer17 jeu17 ven18 3 lun18 jeu18 ven18 sam19 mar19 ven19 sam19 dim20 mer20 sam20 dim lun21 jeu21 dim lun21 mar22 ven22 8 lun22 mar22 mer23 sam23 mar23 mer23 jeu24 Noël dim24 mer24 jeu24 ven25 4 lun25 jeu25 ven25 Vendredi saint sam26 mar26 ven26 DSO sam26 dim27 mer27 sam27 dim lun28 jeu28 dim lun28 Lundi de Pâques mar29 ven29 lun29 mar29 mer30 sam30 mer30 jeu31 dim31 jeu31

40 CERN closure w52 and w53 YETS planning meeting 11/12/1540

41 Access to CERN during end-of-year closure From Saturday 19 December to Sunday 3 January included YETS planning meeting 11/12/1541 Given the current situation, all access to all CERN sites (Meyrin, Prevessin, LHC points etc.) will be strictly controlled during the End-of-Year closure of the organization. Access will be denied, except for those people that are explicitly on a list that will be in the hands of the guards. All groups should therefore provide a list of names of people that would potentially need to enter the CERN premises. In principle this should also contain the names of all non-cern people (outside contractors, users,...) that would need to enter. Any person that would like/need to access and that is not on this list, will have to be authorized by the department head, who will be called for that. See here for more details: https://cern.service-now.com/service- portal/article.do?n=KB https://cern.service-now.com/service- portal/article.do?n=KB

42 P2 ‘ronde’ schedule YETS planning meeting 11/12/1542 Sat19-DecAndre Sun20-DecAndre Mon21-DecAndre Tue22-DecSasha Wed23-DecChilo Thu24-DecChilo Fri25-DecSasha Sat26-DecMartin Sun27-DecMartin Mon28-DecSergey Tue29-DecSergey Wed30-DecArturo Thu31-DecAndre Fri01-JanAndre Sat02-JanArturo Sun03-JanAndre

43 EN-CV: detector cooling status w52-53 YETS planning meeting 11/12/1543 PlantStatus w52-53Contact in case of ALARM SSD/SDDRUNSergey (166783) TPC FECRUNChristian (162430) HMPIDSTBYGiacomo (161834) SPDSTOPSergey (166783) TRDRUN Ana Marin (165596) and Pascal Dillenseger (169441) TOF PHOS EMC DCARUNEMC: Martin (169399) TPC RODRUNChristian:

44 Gas systems status w52-53 YETS planning meeting 11/12/1544 SystemStatusRun supplyBackupComment MCH 61 ST1 ONAr-CO2CO2 MCH 61 ST2 MCH 62 ST3-4-5 MTROFFC2H2F4-iC4H10-SF6 Outlet chambers to atmp. Stop iC4H10 on Dec 14th PMDOFFAr-CO2CO2 CPVOFFAr-CO2N2 HMPIDOFFCH4Flushing by ArgonStop CH4 on Dec 14th HMPID envelopeONN2 manual supply PHOSONN2 manual supply TOFONC2H2F4-SF6C2H2F4 TRDONXe-CO2N2 TPCONAr-CO2 TPC envelopeONCO2 Bronkhorst supply

45 Gas piquet YETS planning meeting 11/12/1545 Andrea D'Auria Frederic Merlet Louis-Philippe De Menezes Louis- Patrick Carrie Steven Pavis Albin Wasem

46 Visits YETS planning meeting 11/12/1546 Underground visits possible in these periods: From Wed 16 th Dec to Fri 18 th Dec included and From Mon 4 th Jan to Thu 3 rd Mar included No visits on Feb 9 th and Feb 26 th

47 AOB YETS planning meeting 11/12/ R-008 (this room) will not be available from 18th January until 2nd February (new Vydio installation)


Laste ned ppt "ALICE 2015-16 YETS Arturo Tauro. Outline YETS planning meeting 11/12/152 YETS duration EN-HE, EN-EL, EN-CV, PH-DT (gas) and GS-ASE activities Acoustic."

Liknende presentasjoner


Annonser fra Google